# Energy shifts and broadening of excitonic resonances in electrostatically-doped semiconductors

### v1

@@ III.3 Optimality and distinguishability of neutral excitons in moiré heterobilayers
III.4 Ruling out screening by resident carriers as a source of decay of excitonic resonances
We have suggested that screening by resident carriers can explain the steady energy redshift of distinct resonances of optimal excitonic complexes. We wish to examine if this screening plays a significant role in the decay and broadening of other optical resonances. For example, the resonances of trions in MoSe2 at $B$==\,=00, marked by $X^{\pm}$ in Fig. 4(b), represent indistinct resonances of optimal complexes. The same is applicable for the positive trion in WSe2, marked by $X^{+}$ in Fig. 4(a). These resonances behave qualitatively differently than distinct resonances of optimal complexes. When the charge density increases, indistinct resonances of optimal complexes decay and broaden, while showing energy blueshift at elevated charge densities. We will first address their decay and broadening and in Section IV we will focus on their energy shift, which goes beyond the considerations of band filling, BGR, and binding energy.
@@ III.4 Ruling out screening by resident carriers as a source of decay of excitonic resonances
Similarly, we can look at dark positive trions whose electron resides in the opposite valley, as shown in Figs. 6(di) and  6(dii). As before, resonances of this optimal trion are said to be distinct or indistinct based on the identity of the recombining hole. Inspecting the resonance $X^{+}_{D,K_{3}}(\sigma_{-})$ in Figs. 6(a) and (b), we see that increasing the hole density renders the phonon-assisted emission narrower and stronger, corroborated by the fact that this is a distinct resonance. On the other hand, the indistinct resonance $X^{+}_{D}(K)$ in Figs. 6(a) and (b) shows significant decay and broadening when the charge density increases. We therefore cannot conclude that screening is causing this decay, because emission from the very same trion is narrower and stronger if it involves recombination of the distinguishable hole.
Figure 7: Helicity resolved magneto-optical reflectance spectra of the neutral exciton in a charge tunable WSe2 monolayer at 4 K. Taken from Ref. [7]. The out-of-plane magnetic field is 17.5 T. As a guide to the eye, the faint dashed lines trace the energy blueshift of the exciton. (b) Helicity resolved optical transitions when the electrons in the monolayer are fully valley polarized. (c) Helicity resolved optical transitions when the holes are fully valley polarized.
III.5 The secondary role of band filling as a source of energy blueshift in absorption
We have identified in Sec. III.1 that distinct resonances of optimal excitonic complexes redshift in energy when resident carriers are added to the system. However, when indistinguishable e-h pairs are photoexcited, absorption resonances trend to blueshift. One should therefore account for the expected energy blueshift from Pauli blocking (i.e., band filling) – namely, the optical energy gap increases because resident carriers occupy the low energy states in the photoexcited valley up to the Fermi energy. Using WSe2 to compare the energy shifts seen in electron and hole doped monolayers, it will become clear that band filling is not the primary reason for the observed energy blueshift of indistinct resonances.
Figure 7(a) shows low-temperature differential reflectance spectra of the neutral exciton in a WSe2 monolayer when $B$=17.5 T [7]. Changing the gate voltage by 1 V in this device amounts to a change of $\sim$4$\times$1011 cm-2 in charge density. Given the large magnetic field, the electrons are fully polarized in the lower CB valley at $-K$ when applying a small positive gate voltage, as shown in Fig. 7(b). Similarly, the holes are fully polarized in the upper VB valley at $K$ when applying a small negative gate voltage, as shown in Fig. 7(c). The gate-voltage window of complete valley polarization is larger for holes because of the larger $g$-factor of the VB valley [11].
We see that the energy blueshift of the exciton is more than 20 meV when the gate voltage increases from 0 to 4 V, and that the blueshift is independent of the helicity (i.e., similar for $\sigma_{\pm}$ polarized light). As we can infer from Fig. 7(b), photoexcitation of either valley can create suboptimal neutral excitons with distinct resonances. The situation is qualitatively different in a fully valley-polarized WSe2 monolayer that is electrostatically-doped with holes. As shown in Fig. 7(c), photoexcitation of the valley at $K$ creates optimal neutral excitons with indistinct resonances. Examining the behavior in Fig. 7(a) when $V_{g}<0$, the energy blueshift of the optimal exciton from $K$ has a staircase shape. As indicated by the optical transition with $\sigma_{+}$ polarization in Fig. 7(c), the staircase shape is governed by progressively blocking the optical transition due to continuous filling of Landau levels with holes. Contrary to this case, the optical transition with $\sigma_{-}$ polarization in Fig. 7(c) means that a suboptimal exciton from $-K$ is not subjected to band filling (i.e., no Pauli blocking). And yet, Fig. 7(a) shows that the energy blueshift of the distinct resonance of the suboptimal exciton from $-K$ is more than 20 meV when the voltage changes from 0 to nearly -4 V whereas the indistinct resonance of the optimal exciton from $K$ experiences similar blueshift when the voltage changes from 0 to -8 V. Namely, the energy blueshift of the neutral exciton is doubled when an empty valley with no Pauli blocking is photoexcited compared with the blueshift that occurs when photoexcitation takes place in an occupied valley (i.e., when Pauli blocking should be strongest). As discussed below, this surprising result can be understood by considering shakeup processes.
@@ IV Shakeup Processes
We have seen that screening is not a primary source of decay, and that band filling is not a major source of energy blueshift in TMD monolayers. These seemingly counterintuitive phenomena can be settled by recognizing that screening by charge carriers and band filling effects are mostly significant when the Fermi energy is comparable to or larger than the binding energy of the e-h pair [50, 58, 59]. In the opposite limit ($\varepsilon_{X}\gg E_{F}$), applicable to TMD systems, we argue that broadening and energy shift of excitonic complexes are instead primarily governed by shakeup processes of resident carriers. In general, there are two types of shakeup processes. The first one involves a breakup (i.e., physical separation) of distinguishable and indistinguishable resident carriers upon photoexcitation (or their reunion after recombination), applicable in the case of indistinct resonances of optimal complexes. The second process involves the evacuation (i.e., removal) of distinguishable resident carriers from the creation region of suboptimal excitonic complexes. We elaborate on these processes in this section.
@@
Figure 8: Colormaps of (a) reflectance and (b) photoluminescence spectra in MoSe2 at $B=0$. Taken from Ref. [23]. The indistinct resonances of the optimal trion complexes, labeled by $X^{\pm}$, broaden when resident carriers are added to the monolayer (readily seen when $|V_{g}|\gtrsim 4$ V). (c) Diagram of the positive trion following photoexcitation of the $K$ valley. (d) Breakup process following the creation of $X^{+}$. (e) Reunion process following the recombination of $X^{+}$. These processes correspond to the broadening and energy blueshift (redshift) in absorption (emission), seen inside the highlighted dashed ovals in (a) and (b).
@@
IV.1 Energy shifts and broadening of indistinct resonances of optimal complexes through breakup and reunion processes
@@
Prior to the absorption of $X^{+}$, resident holes from the VB valleys at $-K$ and $K$ have some spatial overlap. This overlap is illustrated through the wave-packet envelopes of the holes in the left diagram of Fig. 8(d). Mandated by the Pauli exclusion principle, this overlap breaks when the resident hole from $-K$ binds to the photoexcited e-h pair from $K$ to form $X^{+}$, illustrated by the right diagram of Fig. 8(d), where the arrow signifies that resident holes from $K$ are expelled from the region of the created trion. The energy of the absorbed photon is converted to $\hbar\omega=\varepsilon_{T}+\delta_{\varepsilon}$, where $\varepsilon_{T}$ is the trion energy and $\delta_{\varepsilon}$ is the energy required to break the overlap, through which we can understand the energy blueshift. Since the overlap is larger with increasing charge density and since it costs more energy to break a stronger overlap, the energy blueshift is scaled with charge density. Furthermore, the energy blueshift of $X^{+}$ in Fig. 8(a) indicates that $\delta_{\varepsilon}$ is stronger than the underlying redshift effect from screening at elevated densities (Sec. III.2). Finally, the broadening of $X^{+}$ stems from the short lifetime of the perturbed resident-carriers system, and because $\delta_{\varepsilon}$ has a continuum of energy costs depending on whether the breakup involves emission of long wavelength plasmons (with various wavevectors) or trion-hole exchange scattering [18].
@@
Figure 8(e) shows the respective process during recombination. Mandated by the Pauli exclusion principle, the left diagram shows the lack of overlap between the wave-packet envelopes of a hole from $K$ and a trion with indistinguishable e-h pair from $K$. After emission, the overlap between the left-behind hole from $-K$ and a resident hole from $K$ is restored, as shown by their reunion in the right diagram of Fig. 8(e). The recombination process leaves behind an inhomogeneous Fermi sea because the region that was occupied by the trion is left void of indistinguishable resident carriers, which are then pushed back to replenish this region as indicated by the arrow in Fig. 8(e). The energy of the trion before recombination is converted to $\varepsilon_{T}=\hbar\omega+\delta_{\varepsilon}$, where $\delta_{\varepsilon}$ in this case is the energy difference between the inhomogeneous Fermi sea immediately after recombination and the homogeneous one in equilibrium. Since $\delta_{\varepsilon}$ is commensurate with the hole density, the conversion leaves the emitted photon with less energy when the charge density increases, in line with the measured energy redshift inside the highlighted dashed white oval in Fig. 8(b). The redshift is further enhanced by the interplay between BGR and reduced binding energy due to screening (Sec. III.2). As before, the broadening reflects the short lifetime of the perturbed system. That is, the broadening is a measure of the time it takes the inhomogeneous Fermi sea immediately after recombination to restore homogeneity. This response time is commensurate with $\omega_{p}^{-1}$, where $\omega_{p}$ is the plasma frequency.
@@
The self-consistency of this physical picture is examined through the unique case of WSe2 monolayers. Figure 9(a) shows the spectral window of trion emission in a charge tunable WSe2 monolayer at T $=$ 4 K. Details of the sample fabrication and experimental setup can be found in the supplementary information of Ref. [60], and analysis of the dark exciton and trion resonances, $D^{0},\,D^{+},\,D^{-},\,D^{-}_{B}$ can be found in Ref. [18]. Figures 9(b) and (c) show the extracted resonance energies and their full width at half maximum (FWHM), respectively, where the dashed lines show the resonances energies at T $=$ 30 K as well. The numbers in Fig. 9(b) denote the energy redshift of the corresponding resonance in meV per 1011 cm-2. The energy redshift of $H$ and its lack of broadening, as summarized in Figs. 9(b) and (c), are hallmarks of the distinct resonance of the optimal hexciton. The resonance does not broaden since the Fermi sea is not perturbed in this case, and as discussed in Sec. III.2, its energy redshift stems from the interplay between BGR and reduced binding energy due to screening.
@@
The energy redshift of $X^{+}$ in both MoSe2 and WSe2, evidenced from the emission spectra in Figs. 8(b) and  9(a), comes from the fact that their positive trions are optimal complexes with indistinct resonances. Namely, the emission of $X^{+}$ is accompanied by a reunion process. The enhanced broadening of $X^{\pm}$ in Fig. 8(b) compared with that of $X^{+}$ in Fig. 9(a) is attributed to the larger densities used in the MoSe2 device, where --10 V in Fig. 9(a) is the equivalent of --1.5 V in Fig. 8(b). Because recombination of an optimal dark trion at $B=0$ is governed by optical transition of an indistinguishable e-h pair, reunion processes are also the reason that dark trion resonances broaden and redshift in energy, as shown in Fig. 9. Finally, of all resonances in the emission spectra of Figs. 8(b) and  9(a), the only ones that do not shift in energy are the triplet and singlet negative trions ($X^{-}_{S,T}$). As we show next, this behavior is consistent with the fact that $X^{-}_{S,T}$ are distinct rather than indistinct resonances.
@@
Comparing the behavior of $X^{-}_{S,T}$ to that of $H$, we should explain why $X^{-}_{S,T}$ are not subjected to the same screening-induced energy redshift like $H$, and what causes the broadening of $X^{-}_{S,T}$. The broadening indicates a shorter lifetime when the electron density increases, caused by the trion-electron interaction which renders them dark through exchange scattering (i.e., $X^{-}_{S,T}\rightarrow D^{-}$) [18]. The lack of screening-induced energy redshift, on the other hand, is less clear. We can only conjecture that if the resident electrons are quasi-localized when the electrons density $n_{e}$ is a few times $10^{11}$ cm-2 or less, then $X^{-}_{S,T}$ in Fig. 10 are not subjected to the screening-induced energy redshift that $H$ is subjected to when the resident electrons become itinerant at $n_{e}>10^{12}$ cm-2 (Appendix D).
Finally, the emission spectra in Figs. 8(b) and  9(a) show that the energy redshift of indistinct resonances of optimal complexes exist at all charge densities, and that their redshift and broadening are enhanced at elevated charge densities. The enhancement in the case of $X^{\pm}$ in Fig. 8(b) is evident when $|V_{g}|\gtrsim 4\,$$V$, and in the case of $D^{-}$ and $D^{-}_{B}$ in Fig. 9(b) when $V_{g}\gtrsim 15\,$$V$. Our suggested explanation is that reunion processes are relevant regardless of charge density, where the extent of the energy redshift and broadening depend on whether the resident carriers are quasi-localized or itinerant. In more detail, since the Pauli exclusion principle precludes overlap of trions with indistinguishable resident carriers, the rate at which indistinguishable resident carriers creep into the trion region after recombination depends on their itinerancy. The smaller rate of energy redshift and broadening when the resident carriers are quasi-localized is consistent with the fact that Coulomb correlations help indistinguishable and distinguishable resident carriers to avoid each other better at small charge densities [61]. The resulting overlap between indistinguishable and distinguishable resident carriers is smaller than the one expected if they were itinerant, leading to mitigated redshift and broadening effects.
IV.2 Energy blueshift of neutral excitons through evacuation of distinguishable resident carriers
Neutral excitons in TMD monolayers have three possible configurations: a suboptimal exciton with distinct resonance (e.g., Fig. 2(a)), a suboptimal exciton with indistinct resonance (e.g., Fig. 1(a)), or an optimal exciton with indistinct resonance (e.g., Fig. 1(e)). As we discussed in Sec. III.5, while suboptimal excitons with distinct resonances are not subjected to band filling effects (Pauli blocking), they show evidently stronger energy blueshift in absorption experiments compared with the indistinct resonance of the optimal exciton which is subjected to the strongest band filling effect (Fig. 7).
Figure 11: (a) Absorption diagram of a suboptimal neutral exciton with indistinct resonance in electron-doped MoSe2 monolayer. (b)/(c) Corresponding real-space configuration before/after absorption. The bolder arrow of the distinguishable resident carrier indicates a stronger shakeup.
To elucidate the origin of this behavior, we should understand the different roles played by distinguishable and indistinguishable resident carriers. Figure 11(a) shows the photoexcitation of a suboptimal neutral exciton with indistinct resonance, wherein half of the resident electrons are distinguishable ($-K$ valley) and the other half are indistinguishable ($K$ valley). Figures 11(b) and (c) show the conditions before and after the photoexcitation, where the boldness and length of the arrows in Fig. 11(c) reflect the perturbation to distributions of distinguishable and indistinguishable resident carriers. Since the Pauli exclusion principle prohibits the creation of the exciton in the immediate vicinity of indistinguishable resident electrons, it costs more energy to evacuate nearby distinguishable resident carriers with which the e-h pair can bind than to rearrange the more remote indistinguishable resident carriers with which the pair cannot bind [45, 61].

### v2

@@ III.3 Optimality and distinguishability of neutral excitons in moiré heterobilayers
Figure 6: Reflectance spectroscopy of semiconductor quantum wells (QWs). (a) Exciton and trion resonances in CdTe QWs at 1.7 K. Taken from Ref. [50]. (b) The same but in ZnSe QWs at 1.6 K. Taken from Ref. [51]. (c) The same but in GaAs QWs at 2 K. Taken from Ref. [52]. The schemes in the bottom of (a) show zone-center optical transitions from the heavy-hole valence band, where their larger effective mass compared with electrons in the conduction band is illustrated by the valley parabolicity.
III.4 Optimality and distinguishability of excitons and trions in conventional semiconductor QWs
The proposed framework to classify excitonic complexes and their resonances is general and it extends beyond TMD monolayers and their moiré systems. Figure 6 shows resonances of trions and excitons in absorption and reflectivity spectra of various semiconductor quantum wells (QWs) [50, 51, 52]. Figure 6(a) shows the case of CdTe based-QWs at $1.7$ K when the electron density is 1011 cm-2, taken from Kheng et al [50]. At $B=0$, the absorption spectrum includes a single broad resonance that the authors of Ref. [50] marked by FES (Fermi edge singularity). This singularity emerges because of the Coulomb enhancement in 2D systems [67, 68], and its physics will be further discussed in Sec. IV. Based on our proposed classification, we can argue that this resonance is an indistinct resonance of an optimal trion, which one would expect from the optical transition in the left-bottom scheme of Fig. 6(a). Regardless of whether this is a FES resonance or an indistinct resonance of an optimal trion, the reflectance spectrum at $B=11$ T shows a different behavior under the same temperature and electron density. The resonances in this case are attributed to a negative trion with $\sigma_{+}$ polarization (solid line), and a charge-neutral exciton resonance with $\sigma_{-}$ polarization (dashed line). The magnetic field in this case is strong enough to fully polarize the electron gas, rendering the resonance of the optimal trion distinct, as shown by the bottom-middle scheme of Fig. 6(a). Compared with the resonance at $B=0$, the emergence of a narrower trion resonance at lower energy (i.e., redshift) is consistent with the behavior of a distinct resonance of an optimal excitonic complex. Also in agreement with the expected behavior of optimal versus suboptimal excitons, the exciton resonance is observed at $B=11$ T because it becomes optimal, as one can realize from the bottom-right scheme of Fig. 6(a). A similar behavior is observed in TMD monolayers. For example, Figs. 4(a) and (b) show that the resonance of the suboptimal exciton complex ($X^{0}$ at $B=0$) is rapidly suppressed when the charge density increases, whereas Figs. 4(c) and (d) show that $X^{0}$ persists when the exciton becomes optimal at large fields (to be further discussed in Sec. V).
Figure 6(b) shows reflectivity spectra of ZnSe QWs at 1.6 K, taken from Ref. [51]. The bottom and top panels show results at $B=0$ and 7 T, respectively, for three different charge densities. We notice two observations that are self-consistent with our framework. First, the exciton resonance at the larger density is amplified when $B=7$ T (here at $\sigma_{+}$), in accordance with the behavior of an exciton that becomes optimal. Second, the trion resonance (here at $\sigma_{-}$) redshifts in energy without broadening or decay when the charge density increases, in accordance with the behavior of a distinct resonance of an optimal trion (i.e., when its photoexcited e-h pair becomes distinguishable). These trends are similar to the ones seen in TMD monolayers.
Finally, Fig. 6(c) shows the electro-reflectance spectra at progressively increased electron densities in GaAs QWs (bottom to top), taken from Ref. [52]. The temperature is 2 K and the magnetic field is $8$ T. The singlet and triplet resonances, marked by $X^{-}_{s}$ and $X^{-}_{t}$, redshift in energy without broadening when the charge density increases, until they eventually merge at large electron density. Appendix A discusses the emergence of a triplet trion in this QW system. According to our proposed framework, an energy redshift without broadening is to be expected if these trions are optimal and their resonances distinct. This is certainly true for the singlet trion, whose electrons are distinguishable (opposite spins in this case). Identifying the triplet resonance in this case as distinct is more subtle because its electrons have the same spin and both reside in the $\Gamma$-valley of the conduction band in GaAs. Yet, the LL behaves as additional distinguishability knob when the energy difference between LLs exceeds the binding energy of the triplet trion (1 meV in GaAs QWs). The resonance of the triplet trion under a strong magnetic field is distinct because the photoexcited e-h pair weakly binds to an electron in the lowest LL [69], where its large LL degeneracy means that the Pauli principle no longer provides such a strong constraint as at zero field [70]. In agreement with this understanding, the triplet trion is not observed in these QWs at small magnetic fields [69, 70, 71, 72].
@@ III.4 Optimality and distinguishability of excitons and trions in conventional semiconductor QWs
III.5 Ruling out screening by resident carriers as a source of decay of excitonic resonances
We have suggested that screening can explain the steady energy redshift of distinct resonances of optimal excitonic complexes. We wish to examine if this screening plays a significant role in the decay and broadening of other optical resonances. For example, the resonances of trions in MoSe2 at $B$==\,=00, marked by $X^{\pm}$ in Fig. 4(b), represent indistinct resonances of optimal complexes. The same is applicable for the positive trion in WSe2, marked by $X^{+}$ in Fig. 4(a). These resonances behave qualitatively differently than distinct resonances of optimal complexes. When the charge density increases, indistinct resonances of optimal complexes decay and broaden, while showing energy blueshift in absorption at elevated charge densities. We will first address their decay and broadening and in Section IV we will focus on their energy shift, which goes beyond the considerations of band filling, BGR, and binding energy.
@@ III.5 Ruling out screening by resident carriers as a source of decay of excitonic resonances
Similarly, we can look at dark positive trions whose electron resides in the opposite valley, as shown in Figs. 7(di) and  7(dii). As before, resonances of this optimal trion are said to be distinct or indistinct based on the identity of the recombining hole. Inspecting the resonance $X^{+}_{D,K_{3}}(\sigma_{-})$ in Figs. 7(a) and (b), we see that increasing the hole density renders the phonon-assisted emission narrower and stronger, corroborated by the fact that this is a distinct resonance. On the other hand, the indistinct resonance $X^{+}_{D}(K)$ in Figs. 7(a) and (b) shows significant decay and broadening when the charge density increases. We therefore cannot conclude that screening is causing this decay, because emission from the very same trion is narrower and stronger if it involves recombination of the distinguishable hole.
@@ IV Shakeup Processes
We have seen above that screening is not the source of broadening and energy blueshift of excitonic resonances (moreover, in Section V we will also show that band filling is not a primary source of the energy blueshifts observed in TMD monolayers). Here, in this section, we elaborate on the shakeup processes that do cause energy shifts and broadening when the e-h pair is indistinguishable and the excitonic complex is optimal.
The concept of shakeup processes during photoexcitation was introduced originally to understand the Fermi edge problem in metals [79, 75, 76, 77, 78]. In this case, x-ray excitation promotes a deep core electron to the conduction band, leaving behind a hole that remains bound to the atom. The shakeup process is the response of the conduction electrons to the sudden appearance of the localized hole in their vicinity. Around the same time, Mahan predicted a Fermi-edge singularity (FES) in the optical spectra of degenerate semiconductors [80], whose hallmark in the spectrum is an asymmetric resonance that shows energy blueshift and broadening when the charge density increases until the resonance is eventually muted [80, 67, 68]. This prediction was confirmed in semiconductor quantum wells [81], wherein the photoexcited hole was localized by defects or interface fluctuations, rendering its mass ‘infinite’.
To date, investigations that studied the relation between FESs and shakeup processes in degenerate semiconductors [82, 83, 84, 85, 86], have overlooked the key role played by optimality and distinguishability. We have seen that while certain absorption resonances in TMD monolayers show FES-like behavior (blueshift and broadening), the behavior of other resonances in the same monolayer and in the presence of the same Fermi sea is qualitatively different (no blueshift and no broadening). This inconsistency indicates that the understanding of shakeup processes and the conditions at which they emerge have to be refined.
To help us introduce shakeup processes that take into account the optimality and distinguishability, we first explain the relation of these processes to scattering through the lens of Anderson’s infrared catastrophe [87]. Adjusting his model to our problem, we consider $N$ resident electrons in a monolayer and assume that their density is large enough, such that the system is effectively a noninteracting Fermi gas. Upon sudden appearance of a short-range potential (e.g., a photoexcited e-h pair), the overlap between Slater determinants of the $N$-electron systems before and after the perturbation becomes [87]
$S\sim N^{\displaystyle{-\alpha\sin^{2}\delta}},$ (5)
where $\alpha$ is a constant of the order of 1, and $\delta$ is the effective phase shift of Fermi surface electrons because of their scattering with the short-range potential. The infrared catastrophe alluded by Anderson refers to cases where the Slater determinants are orthogonal ($S\rightarrow 0$ when $\delta\neq 0$), meaning that the sudden introduction of the short range potential cannot exist without a shakeup of the Fermi surface electrons. This process involves low-energy excitations of resident electrons to states above the Fermi surface, which can manifest by emission of collective excitations such as plasmons or magnons. On the other hand, a sudden perturbation without this shakeup can be realized if $S\rightarrow 1$ (i.e., $\delta\rightarrow 0$).
@@
Figure 8: Colormaps of (a) reflectance and (b) photoluminescence spectra in MoSe2 at $B=0$. Taken from Ref. [24]. The indistinct resonances of the optimal trion complexes, labeled by $X^{\pm}$, broaden when resident carriers are added to the monolayer (readily seen when $|V_{g}|\gtrsim 4$ V). (c) Diagram of the positive trion following photoexcitation of the $K$ valley. (d) Breakup process following the creation of $X^{+}$. (e) Calculated rates of trion absorption due to a shakeup by plasmons in hole-doped monolayer MoSe2 (solid lines). The calculated rates are normalized with respect to the rate of the direct trion absorption in the low density limit (dotted line). The latter corresponds to absorption without plasmon assistance, where the photon energies are measured with respect to its resonance.
In electrostatically-doped semiconductors, $S\rightarrow 1$ amounts to zero-line optical transitions (i.e., without shakeup of the Fermi sea), described by distinct resonances of optimal excitonic complexes. We explain this scenario through photoexcitation of a negative trion in MoSe2 monolayer whose resident electrons are fully valley polarized, as shown in Fig. 1(f). When the distinguishable e-h pair binds to a resident electron at position $\mathbf{r}$, the Pauli exclusion principle means that no other resident electrons were present at $\mathbf{r}$ and that no other electrons can approach $\mathbf{r}$. Namely, the creation of this trion does not require scattering of the other resident electrons (zero phase shift, $\delta=0$). From a wave-like perspective, the trion inherits the wavevector of the electron and keeps the same crystal momentum after photoexcitation. In other words, the photoexcited e-h pair ‘catches a ride’ with the resident electron to which it binds, so that the Pauli exclusion principle with other resident electrons is not violated. This process describes a different physics than the one provided by the Fermi polaron theory, wherein an attractive polaron is formed when Fermi sea particles approach a photoexcited e-h pair [29, 30]. In the Fermi polaron picture, the resident carriers respond by moving towards the e-h pair, and this motion mandates low-energy excitations across the Fermi surface (i.e., $\delta\neq 0$), which in turn leads to scattering and broadening. However, we have seen that regardless of the semiconductor type, a distinct resonance of an optimal excitonic complex does not broaden when the charge density increases.
Below, we will focus on shakeup processes that accompany the optical transition in the opposite extreme, $S\rightarrow 0$ ($\delta\neq 0$). Applicable to indistinct resonances of optimal complexes, the shakeup involves a breakup (i.e., physical separation) of distinguishable and indistinguishable resident carriers upon photoexcitation, or their reunion after recombination. These processes will help us to self-consistently explain the experimental results.
@@
Energy shifts and broadening of indistinct resonances of optimal complexes through breakup and reunion processes
@@
Prior to the absorption of $X^{+}$, resident holes from the VB valleys at $-K$ and $K$ have some spatial overlap. This partial overlap is illustrated through the wave-packet envelopes of the holes in the top diagram of Fig. 8(d). Mandated by the Pauli exclusion principle, this overlap breaks when the resident hole from $-K$ binds to the photoexcited e-h pair from $K$ to form $X^{+}$, illustrated by the bottom diagram of Fig. 8(d), where the arrow signifies that resident holes from $K$ are forced out of the trion’s region. The energy of the absorbed photon is converted to $\hbar\omega=\varepsilon_{T}+\delta_{\varepsilon}$, where $\varepsilon_{T}$ is the trion energy and $\delta_{\varepsilon}$ is the energy required to break the overlap, through which we can understand the energy blueshift. Since the overlap is larger with increasing charge density and since it costs more energy to break a stronger overlap, the energy blueshift scales with charge density. Furthermore, the energy blueshift of $X^{+}$ in Fig. 8(a) indicates that $\delta_{\varepsilon}$ is stronger than the underlying redshift effect from screening at elevated densities (Sec. III.2). To quantify the spectral broadening due to this shakeup process, we consider the plasmon-assisted absorption of a trion, where plasmon emission creates a charge wave that facilitates the ousting of indistinguishable holes from the trion creation region. Figure 8(e) shows the resulting absorption spectrum, where technical details of this trion-plasmon process are provided in Appendix F. The broadening of $X^{+}$ reflects more than one way to shakeup the system, where photon absorption at a larger energy involves emission of a plasmon with shorter wavelength.
Figure 9: (a) Shakeup (reunion) process following the recombination of $X^{+}$. (b) Calculated rates of trion emission due to a shakeup by plasmons in hole-doped monolayer MoSe2 (solid lines). The calculated rates are normalized with respect to the rate of the direct trion emission in the low density limit (dotted line). The latter corresponds to emission without plasmon assistance, where the photon energies are measured with respect to its resonance.
@@
Figure 9 shows the respective process during recombination. Mandated by the Pauli exclusion principle, the top diagram in Figure 9(a) shows the lack of overlap between the wave-packet envelopes of a hole from $K$ and a trion with indistinguishable e-h pair from $K$. After emission, the overlap between the left-behind hole from $-K$ and a resident hole from $K$ is restored, as shown by their reunion in the bottom diagram of Fig. 9(a). The recombination process leaves behind an inhomogeneous Fermi sea because the region that was occupied by the trion is left void of indistinguishable resident carriers, which are then pushed back to replenish this region as indicated by the arrow in Fig. 9(a). The energy of the trion before recombination is converted to $\varepsilon_{T}=\hbar\omega+\delta_{\varepsilon}$, where $\delta_{\varepsilon}$ in this case is the energy difference between the inhomogeneous Fermi sea immediately after recombination and the homogeneous one in equilibrium. Since $\delta_{\varepsilon}$ is commensurate with the hole density, the conversion leaves the emitted photon with less energy when the charge density increases, in line with the measured energy redshift inside the highlighted dashed white oval in Fig. 8(b). The redshift is further enhanced by the interplay between BGR and reduced binding energy due to screening (Sec. III.2). Figure 9(b) shows the calculated recombination rate when considering trion emission that is accompanied by plasmon emission (Appendix F). The plasmon in this case pushes indistinguishable resident carriers into the region that was left void after recombination, thereby restoring the homogeneity of the Fermi sea.
@@
The self-consistency of the breakup and reunion shakeups is examined through the unique case of WSe2 monolayers. Figure 10(a) shows the spectral window of trion emission in a charge tunable WSe2 monolayer at T $=$ 4 K. Details of the sample fabrication and experimental setup can be found in the supplementary information of Ref. [88], and analysis of the dark exciton and trion resonances, $D^{0},\,D^{+},\,D^{-},\,D^{-}_{B}$ can be found in Ref. [19]. Figures 10(b) and (c) show the extracted resonance energies and their full width at half maximum (FWHM), respectively, where the dashed lines show the resonances energies at T $=$ 30 K as well. The numbers in Fig. 10(b) denote the energy redshift of the corresponding resonance in meV per 1011 cm-2. The energy redshift of $H$ and its lack of broadening, as summarized in Figs. 10(b) and (c), are hallmarks of the distinct resonance of the optimal hexciton. The resonance does not broaden since the Fermi sea is not perturbed in this case, and as discussed in Sec. III.2, its energy redshift stems from the interplay between BGR and reduced binding energy due to screening.
@@
The energy redshift of $X^{+}$ in both MoSe2 and WSe2, evidenced from the emission spectra in Figs. 8(b) and  10(a), comes from the fact that their positive trions are optimal complexes with indistinct resonances. Namely, the emission of $X^{+}$ is accompanied by a reunion process. The enhanced broadening of $X^{\pm}$ in Fig. 8(b) compared with that of $X^{+}$ in Fig. 10(a) is attributed to the larger densities used in the MoSe2 device, where --10 V in Fig. 10(a) is the equivalent of --1.5 V in Fig. 8(b). Because recombination of an optimal dark trion at $B=0$ is governed by optical transition of an indistinguishable e-h pair, reunion processes are also the reason that dark trion resonances broaden and redshift in energy, as shown in Fig. 10. Finally, of all resonances in the emission spectra of Figs. 8(b) and  10(a), the only ones that do not shift in energy are the triplet and singlet negative trions ($X^{-}_{S,T}$). As we show next, this behavior is consistent with the fact that $X^{-}_{S,T}$ are distinct rather than indistinct resonances.
@@
Comparing the behavior of $X^{-}_{S,T}$ to that of $H$, we should explain why $X^{-}_{S,T}$ are not subjected to the same screening-induced energy redshift like $H$, and what causes the broadening of $X^{-}_{S,T}$. The broadening indicates a shorter lifetime when the electron density increases, caused by the trion-electron interaction which renders them dark through exchange scattering (i.e., $X^{-}_{S,T}\rightarrow D^{-}$) [19]. The lack of screening-induced energy redshift, on the other hand, is less clear. We can only conjecture that if the resident electrons are quasi-localized when the electron density $n_{e}$ is a few times $10^{11}$ cm-2 or less, then $X^{-}_{S,T}$ in Fig. 11 are not subjected to the screening-induced energy redshift that $H$ is subjected to when the resident electrons become itinerant at $n_{e}>10^{12}$ cm-2 (Appendix G).
@@
V Energy shift of neutral excitons
We have identified in Sec. III that distinct resonances of optimal excitonic complexes redshift in energy when distinguishable resident carriers are added to the system. In addition, we have identified in Sec. IV that indistinct resonances of optimal excitonic complexes tend to shift and broaden in energy when indistinguishable resident carriers are added to the system. Another interesting phenomenon that we will discuss in this section is that the energy blueshift of the neutral-exciton absorption resonance is relatively weak when the exciton is optimal (i.e., in the absence of distinguishable resident carriers). This result is counterintuitive because adding distinguishable resident carriers does not Pauli block the optical transition, whereas adding indistinguishable resident carriers leads to energy blueshift from the Moss–Burstein effect – namely, the optical energy gap increases when indistinguishable resident carriers occupy the low energy states in the photoexcited valley up to the Fermi energy.